OpenStudy (misssmartiez):

I need help on finding the area of this shape, I must break it apart though. -.-! Fan and medal.

OpenStudy (angel11828):

what shape

OpenStudy (angel11828):

correct

OpenStudy (misssmartiez):

I know that I may do 1/2xb1xh+1/2xb2xh.. However, I don't know what to do first

OpenStudy (misssmartiez):

Really? How?

OpenStudy (gottennis121):

I don't agree.

OpenStudy (angel11828):

9*3= 9*5= + =

OpenStudy (gottennis121):

Here's what I would do: 3 x 9 = ? 9 - 3 x 5 = ? Then add them together

OpenStudy (misssmartiez):

Why do you do that? Sorry confused....

OpenStudy (gottennis121):

The three comes from the top of the figure. You must minus that because you have already done that part of the figure if you first do 3 x 9 to find the area of the tall rectangle to the left.

OpenStudy (angel11828):

Sorry I do have to agree with gottennis121

OpenStudy (misssmartiez):

27 + 6 = 33?...

OpenStudy (gottennis121):

No, 9 x 3 = 27, 6 x 5 = ? Add those.

OpenStudy (gottennis121):

You subtracting 9 - 3, then multiplying by 5

OpenStudy (misssmartiez):

, you said 3 x 5... 3 x 9 = ? 9 - 3 x 5 = ? I did pemdas..

OpenStudy (misssmartiez):

But ok,

OpenStudy (misssmartiez):

57?.

OpenStudy (gottennis121):

Sorry, I think the wrong way, I don't mean to confuse you.

OpenStudy (gottennis121):

Yes! 57 is the correct answer!

OpenStudy (misssmartiez):

Are you sure XD? Because I want a 4/4? (Can you help with another?)

OpenStudy (gottennis121):

it'd be easier if I could talk it out to you. -_-

OpenStudy (misssmartiez):

LOL

OpenStudy (gottennis121):

yes, I'm sure, I took that course.

OpenStudy (gottennis121):

Yes, I can help with one more, then I gotta go...

OpenStudy (gottennis121):

12.5 x 3 = ? Then 3 x 2 = ? Divide by 2

OpenStudy (gottennis121):

I mean divide the 3 x 2, don't include the answer to the once above it.

OpenStudy (misssmartiez):

Really? The lesson said to do this: 1/2xhxb1+1/2xhxb2. I got 21.75, is that right?

OpenStudy (gottennis121):

No. 12.5 x 3 = 37.5 3 x 2 = 5. 5 / 2 = 3.5 Add 37.5 + 3.5 = ?

OpenStudy (misssmartiez):

41?...

OpenStudy (gottennis121):

Hold on, there isn't an answer for that. I'm sure though. You multiply the base x height, then divdie by two for triangle.

OpenStudy (misssmartiez):

oh,

OpenStudy (gottennis121):

3x 2 = 6, sorry. lol

OpenStudy (gottennis121):

basic math mistake.

OpenStudy (gottennis121):

37.5 + 3 = ? There's your answer.

OpenStudy (misssmartiez):

40.5..

OpenStudy (gottennis121):

yes.

OpenStudy (misssmartiez):

Alright, I hope you are correct, I am submitting it.

OpenStudy (misssmartiez):

Oh my gosh thanks! <3 4/4

OpenStudy (gottennis121):

No prob. :)

OpenStudy (misssmartiez):

Yeah. * throws hands up *. I always despised math, but you made it easier.

OpenStudy (gottennis121):

:D Glad I could help.

OpenStudy (misssmartiez):

:3 Cya